OpenStudy (anonymous):

I need help with matrices. Can anyone help explain? equation in reply section.

OpenStudy (anonymous):

\[\left[\begin{matrix}3 & 5 \\ 6 & 2\end{matrix}\right]x=\left[\begin{matrix}-2 & 6 \\ 4 & 12\end{matrix}\right]\] it says to use an inverse matrix to solve each equation or system.

OpenStudy (anonymous):

\[AX=B\]\[X=A ^{-1} B\]

OpenStudy (anonymous):

so i find the inverse of A the multiply that by B?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[\left[\begin{matrix}2 & -5 \\ -6 & 3\end{matrix}\right]\] would be the inverse right?

OpenStudy (anonymous):

where the determinant?

OpenStudy (anonymous):

u should multiply it with\[\frac{ 1 }{ D }\]

OpenStudy (anonymous):

oh right i thought it was 1/ad-bc?

OpenStudy (anonymous):

@ArifL

OpenStudy (anonymous):

yeah that right

OpenStudy (anonymous):

okay i got it thanks so much